The perimeter of a floor is 40 feet and its area is 96 square feet. Find its length and width.
Answer by mananth(16949)

Let length be x and width be y
Perimeter = 40
2x+2y=40
2(x+y)=40
x+y=20
x+y=20
xy=96
..
factors of 96 are 2*2*2*2*2*3
The dimensions are 2*2*3 * 2*2*2
12 by 8 feet